OpenStudy (anonymous):

Simplify the given expression to radical form and justify each step by identifying the properties of rational exponents used. All work must be shown. x to the two–thirds power, over x to the four–ninths power

OpenStudy (anonymous):

First you simplify. Rule of exponents: when dividing you can subtract the exponents if they have the same base. So that gets you to x to the 2/9 power(2/3-4/9=2/9). When you have a fraction as an exponent the top number becomes what you multiply by and the bottom number becomes the root. So in this case the answer would be 2 times the 9th root of x.

OpenStudy (aum):

\(\Large \frac {x^m}{x^n} = x^{m-n}\)

OpenStudy (aum):

\(\Large \frac {x^{2/3}}{x^{4/9}} = x^{2/3-4/9} = x^?\)

OpenStudy (anonymous):

So it would be 2 times the 9th root of x?

OpenStudy (aum):

9th root of x^2 or simply x raised to 2/9. \(\Large x^{2/9} = \sqrt[9]{x^2}\)

OpenStudy (anonymous):

Ohhh ok

OpenStudy (anonymous):

So is that the answer?

OpenStudy (aum):

Yes. I'd probably leave it as \(\Large x^{2/9}\) because the exponents in the original problem are all given as fractions.

OpenStudy (anonymous):

Ok thanks